OpenStudy (anonymous):

What is the slope of the graph of 8x – 4y = 11? a.1/2 b.-1/2 c.-2 d.2

OpenStudy (anonymous):

The standard form of an equation which describes a line is \[y=mx+b\]wherre m is the slope of the line to find the slope bring your equation to the standar form: \[-4y=11-8x=>4y=8x-11=>y=\frac{ 8x-11 }{ 4 }=>y=\frac{ 8x }{ 4 }-\frac{ 11 }{4}\] which is equal to: \[y=2x-\frac{ 11 }{ 4}\]